# Three-body forces and Efimov physics in nuclei and atoms

### v1

@@ 4.1 Overview of the Efimov effect
Among the quantum clusters of few particles, a certain class is remarkable: those which are very close to dissociation into smaller clusters or even all the constituent particles. In quantum systems with short-range interactions, there is indeed a minimum strength of the particles’ attraction that is required for the particles to remain bound to each other. These clusters are thus realised when the attraction strength is just above such critical point, and are therefore relatively weakly bound. What is remarkable about these loosely bound clusters is that they can be very large, much larger than the range of the interactions, thanks to the ability of quantum systems to explore classically forbidden regions. One often speaks of “quantum halos” Jensen2004 when referring to these states, to emphasise their large extent and diluteness. Since a dominant part of their wave function is delocalised outside the region of interaction, it depends upon the interactions only through a few effective parameters. As a result, these states are said to be universal in the sense that only these few parameters are enough to characterise the wave function to good level of approximation, as well as many other properties such as their energy. In other words, different systems with very different interactions can nonetheless lead to the same universal states if their effective interaction parameters are the same.
@@
Among these universal quantum halo states, a more specific class is particularly remarkable and has been a centre of attention of physicists for many years: the so-called “Efimov states”, discovered by V. Efimov in the 1970s Efimov1970a . These three-body clusters occur for inter-particle interactions that are close to the dissociation point of two particles. This means that they can exist either when their two-body subsystems themselves form loosely bound two-body halos ($01/a>$ 0 side of Fig. 11), or even when any of their two-body subsystems is unbound ($01/a<$ 0 side of Fig. 11). In the latter case, these three-body clusters are said to be “Borromean”, a fairly common property of quantum halos. But the unique feature of Efimov states is that they are bound by an effective long-range three-body force called the “Efimov attraction” arising from the short-range two-body forces between the particles. As a result of this long-range attraction, an infinite number of three-body bound states exist before any pair of particles can bind. Moreover, the Efimov attraction is scale invariant, since it decays as $1/R^{2}$, where $R$ is the size of the three-particle system. As a result, the infinity of three-body bound states near the two-body dissociation point is invariant by scaling transformations with scaling factors that are multiples of a certain number $\lambda_{0}$. This property is called “discrete scale invariance” and is depicted in Fig. 11. The figure shows in particular that the energy $E_{n}$ of the $n$th excited state at the two-body dissociation point is related to that of the next one by $E_{n}=\lambda_{0}^{2}E_{n+1}$, which gives the scaling law for the excited states:
@@
Since each three-body bound state is related to the next one by a fixed scaling transformation, it is enough to specify the energy of one particular three-body bound state to determine the energies of all other bound states. This single energy scale fixing the whole spectrum is referred to as the “three-body parameter” and is conventionally defined to be the limit of $E_{n}\lambda_{0}^{2n}$ for large $n$.
Figure 11: Efimov plot: schematic plot of three-particle energy $E$ as a function of the inverse of the scattering length $a$ between particles, which is measure of the particles’ attraction strength indicated by the wide purple arrow. The energy $E$ is rescaled into a wave number $k=\text{sign}(E)\sqrt{m|E|}/\hbar$ (where $m$ is the mass of the particles and $\hbar$ the reduced Planck constant) so that both coordinates are homogeneous to an inverse length, and can be expressed in units of the inverse of the range $b$ of the particles’ interactions. As the attraction strength is increased, the cluster of two particles (black line) appears from the point $01/a=$ 0 indicated by the black arrow. Conversely, this point can also be regarded as the two-body dissociation point when the attraction strength is decreased. There is an infinite number of three-particle clusters appearing before that point. Their energies form a discrete-scale invariant pattern, whereby any point on these energy curves can be mapped to another point by scaling the coordinates ($k,1/a$) by a factor $\lambda_{0}$, as indicated by the red arrows.
The occurrence of discrete-scale invariance in systems of particles with short-range interactions is called the “Efimov effect”. There are certain conditions for the Efimov effect to occur. First of all, the quantum statistics and spin of the particles play an important role, because the Pauli exclusion directly competes with the Efimov attraction efimov1973energy ; PhysRevA.67.010703 ; kartavtsev2007low ; PhysRevLett.103.153202 ; PhysRevA.86.062703 ; endo2011universal . While identical bosons are always subject to the Efimov effect whenever they are close to their two-body dissociation point, identical fermions with spin 1/2 or less (polarised fermions) cannot exhibit the Efimov effect due to the Pauli exclusion. Generally, for the Efimov effect to occur, at least two pairs of particles must be able to interact in the $s$-wave, close to their dissociation point. This means that their corresponding $s$-wave scattering lengths must be, in absolute value, much larger (typically 10 times PhysRevLett.107.120401 ; PhysRevLett.108.263001 ; pascaleno3BP1 ; pascaleno3BP2 ; johansen2017testing ; PhysRevLett.111.053202 ) than the range of their interactions. This is a rather stringent requirement, since most systems have scattering lengths of the order of the interaction range, and therefore do not exhibit the Efimov effect.
@@ 4.3 Efimov states in nuclear physics
Systems of two neutrons interacting with a nucleus are more promising candidates for Efimov states. Although a single neutron tends to be absorbed by a nucleus due to the strong force, the situation changes when the number of neutrons becomes large and approaches the neutron dripline where neutrons is no longer tighty bound around the nucleus. Around this dripline, it may happen for certain nuclides that they could almost but not quite bind an extra neutron. If one regards the nucleus and the neutron as two distinct particles, this situation corresponds to the two-body dissociation point discussed earlier, around which the Efimov effect can occur. It is thus expected that, in the presence of two neutrons, such a nucleus could form a three-body Efimov state, where the two neutrons would remain at large distance forming a halo around the nucleus. Indeed, nuclei with two-neutron halo structures have been studied both experimentally and theoretically since the 1980s, such as 6He, 11Li, 12Be, 17B, 19B, 20C, 22C, 62Ca, and 72Ca tanihata1985measurements ; AlKhalili2004 ; Tanihata2013 ; Fedorov1994 ; Amorim1997 ; AnnRev_HamPlatt ; naidon2017efimov ; FREDERICO2012939 ; Canham2008 ; Hammer_2017 ; ACHARYA2013196 ; PhysRevLett.111.132501 ; PhysRevC.105.024310 ; PhysRevLett.120.052502 . Compared to the triton, these systems stand as better candidates for Efimov states Fedorov1994 ; Amorim1997 ; Canham2008 because some of them would correspond to the Borromean regime (left-hand side of Fig. 11) where the universal description in terms of Efimov state is more effective. It is indeed widely believed that these two-neutron halo nuclei are examples of Efimov states. Three-body models of some nuclei seem to reproduce observations and confirm this view.
@@
Interestingly, the geometry of Efimov states near three-body dissociation tends to the universal halo geometry mentioned in Sec. 4.2. Thus, if two-neutron halo nuclei are indeed Borromean Efimov states, they should approach this halo universality. However, the neutron-neutron scattering length is somewhat too small with respect to $b$ to fully reach the universal halo regime, so that they only constitute approximations of such halos. Among the known two-neutron halo nuclei, 19B and 22C are good candidates to exhibit the halo universality naidon2023universal . This could be experimentally confirmed by improving the measurements of their various mean square radii and binding energies.
@@ 4.4.1 Helium-4
Atomic physics has been more favourable for the demonstration of the Efimov effect. Soon after its prediction, it was realised that helium-4 atoms are good candidates for the observation of Efimov states, since they are bosonic and their $s$-wave scattering length happens to be about 20 times larger than their interaction range. This allows the existence of two trimer states, a ground state and an excited state Lim1977 . There was a long experimental effort to observe these states, which involved diffracting beams of helium clusters through a grating, selecting the deflecting beam corresponding to trimers and analysing their geometry. Although the ground state was observed from the 1990s Schoellkopf1994 ; Bruehl2005 , it does not fully conform to the Efimov zero-range theory. Observation of the excited state was finally reported in 2015 Kunitski2015 , thanks to the Coulomb explosion imaging technique. It revealed that its geometry excellently conform to the Efimov theory, although the universal scaling ratio $\lambda_{0}$ could not be directly confirmed.
@@ 4.4.2 Cold atoms
The simplest way to reveal the presence of Efimov states in this range is to monitor the losses occurring among atoms through three-body collisions and recombinations into deeply bound dimers PhysRevLett.83.1751 ; PhysRevLett.85.908 . Indeed, nearby the scattering lengths at which an Efimov state appears at the three-body zero-energy threshold, these losses are significantly enhanced by the three-body resonances above threshold associated to these Efimov states Kraemer2006 ; PhysRevLett.103.163202 ; zaccanti2009observation ; pollack2009universality . Since these scattering lengths are related to each other by a scaling with the universal scaling ratio $\lambda_{0}$, it makes possible the experimental determination of this ratio. The discrete scale invariance was thus confirmed in both systems of identical bosons Huang2014 and systems of heavy bosons interacting with a lighter boson Tung2014 ; Pires2014 . The measured scaling ratios were found to be consistent with the universal predictions, within the experimental uncertainties and corrections due to the finite-range of interactions. To minimize the effects of finite temperatures, trap size, and other aspects deteriorating Efimov signatures, there is an on-going attempt to observe Efimov states in the microgravity environment aboard the International Space Station (ISS) elliott2018nasa ; oudrhiri2023nasa ; frye2021bose .
@@
More recently, an ultracold mixture of Er and Li atoms has been realized PhysRevA.107.L031306 . The Er and Li atoms are cooled down by a sympathetic cooling method utilizing the Yb atoms as a coolant, which is so powerful that a new kind of quantum degenerate gases of dual BECs of highly-magnetic atoms of Er and totally-nonmagnetic atoms of Yb can also successfully be created. The Er-Li system (mass ratio $\approx 28$) is deemed as a major candidate to realize the fermionic Efimov states. As both Er and Li atoms have open-shell atomic structures, they may have magnetic Feshbach resonances broad enough to accurately control the scattering length PhysRevA.92.022708 . Recently, systematic measurements of Er-Li Feshbach resonances for various isotope mixtures have been performed in a magnetic field range below 1 kG, and succeeded in observing many Feshbach resonances schafer_feshbach_2022 ; 167Er6Li2023 . While the observed Feshbach resonances are mostly narrow, possibly induced by the anisotropic electrostatic interactions or dipolar interactions chen2015anisotropy , several broad Feshbach resonances were also found
The observed broad Feshbach resonances in Er-Li is an important step toward observing the fermionic Efimov trimers. The Efimov trimers are expected to be observed from the loss measurement. In particular, the loss process involving the Er-Er-Li channel, namely two Er accompanied by one Li atoms should be predominantly enhanced. In Ref 167Er6Li2023 , the atom loss behavior around the Feshbach resonance at $455$ G is observed in detail for the 167Er and 6Li mixture. The atom loss is found to be predominantly induced through the Er-Er-Li channel, namely two Er accompanied by one Li, which is consistent with the formation of the fermionic Efimov trimers. It is also shown that the shape of the observed resonance is asymmetric, which is also consistent with the previous observations of multiple losses originated from the formation of several Efimov trimers Pires2014 ; Tung2014 . More detailed measurements, including the determination of the inter-species scattering lengths, will reveal the behaviors of these resonant losses induced by the fermionic Efimov trimers in this mass-imbalanced Er-Li mixture. Due to the strong magnetic dipole interaction between Er atoms, an interplay of the van der Waals and dipole interaction is predicted to lead to novel universal behaviors of the Efimov states OiEndo2024 . We also note that the Er-Li mixture is also a fascinating candidate for realizing a novel $p$-wave superfluid of fermions of Er through the induced attraction via a BEC of light Li atoms PhysRevA.61.053601 ; PhysRevLett.117.245302 ; gurarie2007resonantly ; yamaguchi2017novel , which is also interesting from the viewpoint of the effect of the many-body background of the novel Efimov trimers.
@@ Acknowledgements.
This work is supported by Grant-In-Aid for Scientific Research on Innovative Areas ‘Clustering as a window on the hierarchical structure of quantum systems’. S.E. is supported by JSPS KAKENHI Grant Nos. JP21H00116, 22K03492, and 23H01174. S.E. acknowledges support from Institute for Advanced Science, University of Electro-Communications. E.E. is supported by ERC under the EU Horizon 2020 research and innovation programme (ERC AdG Nuclear Theory, No. 885150), by DFG and NSFC through funds provided to the Sino-German CRC 110 “Symmetries and the Emergence of Structure in QCD” (DFG Project ID 196253076 - TRR 110, NSFC Grant No. 11621131001), by the MKW NRW (funding code NW21-024-A), and by the EU Horizon 2020 research and innovation programme (STRONG-2020, No. 824093). E.E. acknowledges the work of his collaborators Ashot Gasparyan, Hermann Krebs, Ulf-G. Meißner, and the members of the LENPIC Collaboration. P.N. is supported by JSPS KAKENHI Grant No. JP23K03292. Y.N. is supported by JSPS KAKENHI Grant Nos. JP18H05405 and JP21K03384. K.S. is supported by JSPS KAKENHI Grant Nos. JP25105502, JP16H02171, JP18H05404, JP20H05636, and JST ERATO Grant No. JPMJER2304, Japan. Y.T. is supported by the Grant-in-Aid for Scientific Research of JSPS (Nos. JP17H06138, JP18H05405, JP18H05228, JP21H01014, JP21K03384, and JP22K20356), JST CREST (Nos. JPMJCR1673 and JPMJCR23I3), MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant No. JPMXS0118069021, and JST Moonshot R&D Grant No. JPMJMS2269.

### v2

@@ 4.1 Overview of the Efimov effect
Among the quantum clusters of few particles, a certain class is remarkable: those which are very close to dissociation into smaller clusters or even all the constituent particles. In quantum systems with short-range interactions, there is indeed a minimum strength of the particles’ attraction that is required for the particles to remain bound to each other. These clusters are thus realised when the attraction strength is just above such critical point, and are therefore relatively weakly bound. What is remarkable about these loosely bound clusters is that they can be very large, much larger than the range of the interactions, thanks to the ability of quantum systems to explore classically forbidden regions. One often speaks of “quantum halos” Jensen2004 when referring to these states, to emphasise their large extent and diluteness. Since a dominant part of their wave function is delocalised outside the region of interaction, it depends upon the interactions only through a few effective parameters. As a result, these states are said to be universal in the sense that only these few parameters are enough to characterise the wave function to good level of approximation, as well as many other properties such as their energy. In other words, different systems with very different interactions can nonetheless lead to the same universal states if their effective interaction parameters are the same. These universal parameters correspond to the low-energy constants in the ChEFT discussed in Sec. 2.2; The leading order constant arising from the low-energy two-body interaction, the scattering length $a$, is the prominent one.
@@
Among these universal quantum halo states, a more specific class is particularly remarkable and has been a centre of attention of physicists for many years: the so-called “Efimov states”, discovered by V. Efimov in the 1970s Efimov1970a . These three-body clusters occur for inter-particle interactions that are close to the dissociation point of two particles. This means that they can exist either when their two-body subsystems themselves form loosely bound two-body halos ($01/a>$ 0 side of Fig. 11), or even when any of their two-body subsystems is unbound ($01/a<$ 0 side of Fig. 11). In the latter case, these three-body clusters are said to be “Borromean”, a fairly common property of quantum halos. But the unique feature of Efimov states is that they are bound by an effective long-range three-body force called the “Efimov attraction” arising from the short-range two-body forces between the particles. As a result of this long-range attraction, an infinite number of three-body bound states exist before any pair of particles can bind. Moreover, the Efimov attraction is scale invariant, since it decays as $1/R^{2}$, where $R$ is the size of the three-particle system. As a result, the infinity of three-body bound states near the two-body dissociation point is invariant by scaling transformations with scaling factors that are multiples of a certain number $\lambda_{0}$. This property is called “discrete scale invariance” and is depicted in Fig. 11. The figure shows in particular that the energy $E_{n}$ of the $n$-th excited state at the two-body dissociation point is related to that of the next one by $E_{n}=\lambda_{0}^{2}E_{n+1}$, which gives the scaling law for the excited states:
@@
Since each three-body bound state is related to the next one by a fixed scaling transformation, it is enough to specify the energy of one particular three-body bound state to determine the energies of all other bound states. This single energy scale fixing the whole spectrum is referred to as the “three-body parameter” and is conventionally defined to be the limit of $E_{n}\lambda_{0}^{2n}$ for large $n$. The three-body parameter corresponds to the dominant low-energy constant involving three particles (see N2LO in Fig. 3). Along with the scattering length, these are the two universal parameters that characterise the Efimov states.
Figure 11: Efimov plot: schematic plot of three-particle energy $E$ as a function of the inverse of the scattering length $a$ between particles, which is a measure of the particles’ attraction strength indicated by the wide purple arrow. The energy $E$ is rescaled into a wave number $k=\text{sign}(E)\sqrt{m|E|}/\hbar$ (where $m$ is the mass of the particles and $\hbar$ the reduced Planck constant) so that both coordinates are homogeneous to an inverse length, and can be expressed in units of the inverse of the range $b$ of the particles’ interactions. As the attraction strength is increased, the cluster of two particles (black line) appears from the point $01/a=$ 0 indicated by the black arrow. Conversely, this point can also be regarded as the two-body dissociation point when the attraction strength is decreased. There is an infinite number of three-particle clusters appearing before that point. Their energies form a discrete-scale invariant pattern, whereby any point on these energy curves can be mapped to another point by scaling the coordinates ($k,1/a$) by a factor $\lambda_{0}$, as indicated by the red arrows.
The occurrence of discrete-scale invariance in systems of particles with short-range interactions is called the “Efimov effect”. There are certain conditions for the Efimov effect to occur. First of all, the quantum statistics and spin of the particles play an important role, because the Pauli exclusion directly competes with the Efimov attraction efimov1973energy ; PhysRevA.67.010703 ; kartavtsev2007low ; PhysRevLett.103.153202 ; PhysRevA.86.062703 ; endo2011universal . While identical bosons are always subject to the Efimov effect whenever they are close to their two-body dissociation point, identical fermions with spin 1/2 or less (polarised fermions) cannot exhibit the Efimov effect due to the Pauli exclusion. Generally, for the Efimov effect to occur, at least two pairs of particles must be able to interact in the $s$ wave, close to their dissociation point. This means that their corresponding $s$-wave scattering lengths must be, in absolute value, much larger (typically 10 times PhysRevLett.107.120401 ; PhysRevLett.108.263001 ; pascaleno3BP1 ; pascaleno3BP2 ; johansen2017testing ; PhysRevLett.111.053202 ) than the range of their interactions. This is a rather stringent requirement, since most systems have scattering lengths of the order of the interaction range, and therefore do not exhibit the Efimov effect.
@@ 4.3 Efimov states in nuclear physics
Systems of two neutrons interacting with a nucleus are more promising candidates for Efimov states. Although a single neutron tends to be absorbed by a nucleus due to the strong force, the situation changes when the number of neutrons becomes large and approaches the neutron dripline where neutrons are no longer tightly bound around the nucleus. Around this dripline, it may happen for certain nuclides that they could almost but not quite bind an extra neutron. If one regards the nucleus and the neutron as two distinct particles, this situation corresponds to the two-body dissociation point discussed earlier, around which the Efimov effect can occur. It is thus expected that, in the presence of two neutrons, such a nucleus could form a three-body Efimov state, where the two neutrons would remain at large distance forming a halo around the nucleus. Indeed, nuclei with two-neutron halo structures have been studied both experimentally and theoretically since the 1980s, such as 6He, 11Li, 12Be, 17B, 19B, 20C, 22C, 62Ca, and 72Ca tanihata1985measurements ; AlKhalili2004 ; Tanihata2013 ; Fedorov1994 ; Amorim1997 ; AnnRev_HamPlatt ; naidon2017efimov ; FREDERICO2012939 ; Canham2008 ; Hammer_2017 ; ACHARYA2013196 ; PhysRevLett.111.132501 ; PhysRevC.105.024310 ; PhysRevLett.120.052502 . Compared to the triton, these systems stand as better candidates for Efimov states Fedorov1994 ; Amorim1997 ; Canham2008 because some of them would correspond to the Borromean regime (left-hand side of Fig. 11) where the universal description in terms of Efimov state is more effective. It is indeed widely believed that these two-neutron halo nuclei are examples of Efimov states. Three-body models of some nuclei seem to reproduce observations and confirm this view.
@@
Interestingly, the geometry of Efimov states near three-body dissociation tends to the universal halo geometry mentioned in Sec. 4.2. Thus, if two-neutron halo nuclei are indeed Borromean Efimov states, they should approach this halo universality. However, the neutron-neutron scattering length is somewhat too small with respect to $b$ to fully reach the universal halo regime, so that they only constitute approximations of such halos. Among the known two-neutron halo nuclei, 19B and 22C are good candidates to exhibit the halo universality naidon2023universal . It can be tested through enhanced measurements of the mean square radii and binding energies, or more directly by measuring the two-neutron correlation PhysRevLett.125.252501 .
@@ 4.4 Efimov states in atomic physics
Atomic physics has been more favourable for the demonstration of the Efimov effect. Below we review experimental progress in observing the Efimov states in atomic systems.
@@ 4.4.1 Helium-4
Soon after its prediction, it was realised that helium-4 atoms are good candidates for the observation of Efimov states, since they are bosonic and their $s$-wave scattering length happens to be about 20 times larger than their interaction range. This allows the existence of two trimer states, a ground state and an excited state Lim1977 . There was a long experimental effort to observe these states, which involved diffracting beams of helium clusters through a grating, selecting the deflecting beam corresponding to trimers and analysing their geometry. Although the ground state was observed from the 1990s Schoellkopf1994 ; Bruehl2005 , it does not fully conform to the Efimov zero-range theory. Observation of the excited state was finally reported in 2015 Kunitski2015 , thanks to the Coulomb explosion imaging technique. It revealed that its geometry excellently conforms to the Efimov theory, although the universal scaling ratio $\lambda_{0}$ could not be directly confirmed.
@@ 4.4.2 Cold atoms
The simplest way to reveal the presence of Efimov states is to monitor the losses occurring among atoms through three-body collisions and recombinations into deeply bound dimers PhysRevLett.83.1751 ; PhysRevLett.85.908 . Indeed, nearby the scattering lengths at which an Efimov state appears at the three-body zero-energy threshold, these losses are significantly enhanced by the three-body resonances above threshold associated to these Efimov states Kraemer2006 ; PhysRevLett.103.163202 ; zaccanti2009observation ; pollack2009universality . Since these scattering lengths are related to each other by a scaling with the universal scaling ratio $\lambda_{0}$, it makes possible the experimental determination of this ratio. The discrete scale invariance was thus confirmed in both systems of identical bosons Huang2014 and systems of heavy bosons interacting with a lighter boson Tung2014 ; Pires2014 . The measured scaling ratios were found to be consistent with the universal predictions, within the experimental uncertainties and corrections due to the finite-range of interactions. To minimize the effects of finite temperatures, trap size, and other aspects deteriorating Efimov signatures, there is an on-going attempt to observe Efimov states in the microgravity environment aboard the International Space Station (ISS) elliott2018nasa ; oudrhiri2023nasa ; frye2021bose .
@@
More recently, an ultracold mixture of Er and Li atoms has been realized PhysRevA.107.L031306 . The Er and Li atoms are cooled down by a sympathetic cooling method utilizing the Yb atoms as a coolant, which is so powerful that a new kind of quantum degenerate gases of dual BECs of highly-magnetic atoms of Er and totally-nonmagnetic atoms of Yb can also successfully be created. The Er-Li system (mass ratio $\approx 28$) is deemed as a major candidate to realize the fermionic Efimov states. As both Er and Li atoms have open-shell atomic structures, they may have magnetic Feshbach resonances broad enough to accurately control the scattering length PhysRevA.92.022708 . Recently, systematic measurements of Er-Li Feshbach resonances for various isotope mixtures have been performed in a magnetic field range below 1 kG, and succeeded in observing many Feshbach resonances schafer_feshbach_2022 ; 167Er6Li2023 . While the observed Feshbach resonances are mostly narrow, possibly induced by the anisotropic electrostatic interactions or dipolar interactions chen2015anisotropy , several broad Feshbach resonances were also found.
The observed broad Feshbach resonances in Er-Li is an important step toward observing the fermionic Efimov trimers. The Efimov trimers are expected to be observed from loss measurement. In particular, the loss process involving the Er-Er-Li channel, namely two Er accompanied by one Li atoms should be predominantly enhanced. In Ref. 167Er6Li2023 , the atom loss behavior around the Feshbach resonance at $455$ G is observed in detail for the 167Er and 6Li mixture. The atom loss is found to be predominantly induced through the Er-Er-Li channel, namely two Er accompanied by one Li, which is consistent with the formation of the fermionic Efimov trimers. It is also shown that the shape of the observed resonance is asymmetric, which is also consistent with the previous observations of multiple losses originating from the formation of several Efimov trimers Pires2014 ; Tung2014 . More detailed measurements, including the determination of the inter-species scattering lengths, will reveal the behaviors of these resonant losses induced by the fermionic Efimov trimers in this mass-imbalanced Er-Li mixture. Due to the strong magnetic dipole interaction between Er atoms, an interplay of the van der Waals and dipole interaction is predicted to lead to novel universal behaviors of the Efimov states OiEndo2024 . We also note that the Er-Li mixture is also a candidate for realizing a novel $p$-wave superfluid of fermions of Er through the induced attraction via a BEC of light Li atoms PhysRevA.61.053601 ; PhysRevLett.117.245302 ; gurarie2007resonantly ; yamaguchi2017novel , which is also interesting from the viewpoint of the effect of the many-body background of the novel Efimov trimers.
@@ Acknowledgements.
This work is supported by Grant-In-Aid for Scientific Research on Innovative Areas ‘Clustering as a window on the hierarchical structure of quantum systems’. S.E. is supported by JSPS KAKENHI Grant Nos. JP21H00116, 22K03492, 23H01174, and Matsuo Foundation. S.E. acknowledges support from Institute for Advanced Science, University of Electro-Communications. E.E. is supported by ERC under the EU Horizon 2020 research and innovation programme (ERC AdG Nuclear Theory, No. 885150), by DFG and NSFC through funds provided to the Sino-German CRC 110 “Symmetries and the Emergence of Structure in QCD” (DFG Project ID 196253076 - TRR 110, NSFC Grant No. 11621131001), by the MKW NRW (funding code NW21-024-A), and by the EU Horizon 2020 research and innovation programme (STRONG-2020, No. 824093). E.E. acknowledges the work of his collaborators Ashot Gasparyan, Hermann Krebs, Ulf-G. Meißner, and the members of the LENPIC Collaboration. P.N. is supported by JSPS KAKENHI Grant No. JP23K03292. Y.N. is supported by JSPS KAKENHI Grant Nos. JP18H05405 and JP21K03384. K.S. is supported by JSPS KAKENHI Grant Nos. JP25105502, JP16H02171, JP18H05404, JP20H05636, and JST ERATO Grant No. JPMJER2304, Japan. Y.T. is supported by the Grant-in-Aid for Scientific Research of JSPS (Nos. JP17H06138, JP18H05405, JP18H05228, JP21H01014, JP21K03384, and JP22K20356), JST CREST (Nos. JPMJCR1673 and JPMJCR23I3), MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant No. JPMXS0118069021, and JST Moonshot R&D Grant No. JPMJMS2269.